OpenStudy (dariel7):

The snack bar at your school has added sushi to its menu. The ingredients for one roll includes sushi rice,seaweeds and 4 oz of smoke salmon. One roll can be cut into 9 servings. Write an expression for the amount salmon needed to make x servings of sushi. How much salmon is needed to make 18 servings? 27 servings? 90 servings?

OpenStudy (unklerhaukus):

4 oz = 1 roll = 9 servings

OpenStudy (dariel7):

Answer should be in algebraic expression.

OpenStudy (unklerhaukus):

9 s = 4 oz 1 s = 4/9 oz x s = 18 s = 18 * 4/9 oz 27 s = 90 s =